OpenStudy (anonymous):

Using the least integer principle, define r to be least integer for which j - qk is positive. Prove that 0< r <=k

OpenStudy (zzr0ck3r):

? this shuold that q or j be an r?

OpenStudy (anonymous):

Actually r= j - qk. This comes form Euclid's division lemma: j=qk + r

OpenStudy (thomas5267):

Is j and q positive or not?

OpenStudy (thomas5267):

Proofwiki has a rigorous but rather hard to understand proof. Ask if you need help. https://proofwiki.org/wiki/Division_Theorem

OpenStudy (anonymous):

I think I understood the informal proof. Just to make sure, when it says that: b∖(r−r0). −b<r−r0<b. Hence, r−r0=0 r-r0 must be 0 because there is no other way it is divisible by b. Am I right?